2 LN 2 - ((LN 36) / 2) + (2/3 LN 27). I think it is: LN [(27 ^ (2/3))(2^2) / (36^.5) Is this how I rewrite it.
Found 2 solutions by stanbon, 12cleo:
Answer by stanbon(75887)

2 LN 2 - ((LN 36) / 2) + (2/3 LN 27).
I think it is: LN [(27 ^ (2/3))(2^2) / (36^.5) Is this how I rewrite it.
=========================================================
2 LN(2) - ((LN(36) / 2) + (2/3) LN 27
-------------------
= ln(4) - ln(36^(1/2)) + ln(27^(2/3))
-------------------
= ln(4) - ln(6) + ln(9)
------
= ln[4*9/6]
------
= ln(6)
=================
Cheers,
Stan H.

Answer by 12cleo(1)